OpenStudy (anonymous):

Let f(x)=-2x-4 and g(x)=-7x-5. Find f x g and its domain

OpenStudy (anonymous):

\[fxg(x)=f(g(x))=f(-7x-5)=-2(-7x-5)-4=14x+10-4=14x+6\] Domain for fxg(x) is set of all real numbers.\[\mathbb{R}\]